SPECTROSCOPY OF THE CH_{3}-HCL COMPLEX IN HELIUM NANODROPLETS

\underline{CHRISTOPHER P. MORADI} and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556.

The CH_{3}-HCl van der Waals complex has been isolated in superfluid helium nanodroplets, and the infrared-active hydrogen stretching vibrations have been identified. A 153 cm^{−1}redshift of the HCl stretch is observed (relative to the HCl monomer peak at 2905 cm^{−1}), consistent with the predicted hydrogen bonding interaction between the two molecules. The perpendicular CH stretching transitions originating from (N,K ) = (0,0) and (1,1) are also observed, and simulations indicate that the band origin is shifted by roughly 6 cm^{−1}relative to the band origin of methyl radical in helium droplets^{a} (3160.9(1) cm^{−1}). The barrier height to formation of CH_{4} + Cl from CH_{3} + HCl is only 1.5 kcal mol^{−1} with a 0 K reaction exothermicity of about 1.2 kcal mol^{−1}.^{b} The broad linewidth of the parallel transition, relative to the perpendicular transition, indicates that excitation of the HCl stretch may be lifetime broadened due to optical pumping of the weakly bound complex over the barrier to form CH_{4} + Cl. Additionally, one of the HCl stretches due to CH_{3}-(HCl)_{2} has been identified and lies approximately 27 cm^{−1}to the red of the “bound” stretch of the HCl dimer at 2852 cm^{−1}. The feasibility of an IR-IR double resonance experiment to probe for the photo-induced formation of CH_{4}-Cl will be discussed.

aA. M. Morrison, P. L. Raston, G. E. Douberly. J. Phys. Chem. A. 2013, in press.

bA. J. Orr-Ewing, C. Murray. Int. Rev. Phys. Chem. 2004, 4, 435-482.